# Challenges and Opportunities for Visual Analytics in Jurisprudence

### v2

@@ Abstract
Legal exploration, analysis, and interpretation remain complex and demanding tasks, even for experienced legal scholars, due to the domain-specific language, tacit legal concepts, and intentional ambiguities embedded in legal texts. In related, text-based domains, Visual Analytics (VA) and Large Language Models (LLMs) have become indispensable tools for navigating documents, representing knowledge, and supporting analytical reasoning. However, legal scholarship presents distinct challenges: it requires managing formal legal structure, drawing on tacit domain knowledge, and documenting intricate and accurate reasoning processes – needs that current VA systems designs and LLMs fail to address adequately. We identify previously unexamined key challenges and underexplored opportunities in applying VA to jurisprudence to explore how these technologies might better serve the legal domain. Based on semi-structured interviews with nine legal experts, we find a significant gap in tools and means that can externalize tacit legal knowledge in a form that is both explicit and machine-interpretable. Hence, we propose leveraging interactive visualization for this articulation, teaching the machine relevant semantic relationships between legal documents that inform the predictions of LLMs, facilitating the enhanced navigation between hierarchies of legal collections. This work introduces a user-centered VA workflow to the jurisprudential context, recognizing tacit legal knowledge and expert experience as vital components in deriving legal insight, comparing it with established practices in other text-based domains, and outlining a research agenda that offers future guidance for researchers in Visual Analytics for law and beyond.
@@
Visual Analytics, Large Language Models, Tacit Knowledge, Jurisprudence
@@ 1 Introduction
Legal text data is vast [1, 2, 3] but inherently structured [4, pp. 92–93]. It combines domain-specific language [4, p. 123] and precise phraseology with intentional ambiguity at times [4, p. 128]. This ambiguity allows for statutory interpretation [4, p. 22], enabling laws to evolve alongside societal norms and values. For instance, in 2018, the German Federal Constitutional Court interpreted the protection of life and physical integrity according to Art. 2 Abs. 2 Satz 1 GG to include protection against the dangers of climate change [5]. The legal text merely states that “every person shall have the right to life and physical integrity. Freedom of the person shall be inviolable.” This interpretation underscores that considering legal norms in isolation is insufficient; comprehensive statutory interpretation also relies on an extensive understanding of fundamental and specific legal principles [4, pp. 92–93, 460].
@@
These legal principles are formed and conceived within individual jurisdictions and a legal system. Several legal systems co-exist in our world, where common law and civil law are predominant [6], the latter being the most widespread. While the Anglo-American common law system relies on judicial precedents and case law, the civil law system, rooted in codes and statutes, lends itself to more structured analysis. These differences are the focus of comparative law studies, reaching back centuries [7, 8]. Seminal work [9, 10] has analyzed how methodologies and approaches are comparable. Nevertheless, legal practices can differ between jurisdictions, even within the same legal system. In particular, with more tight integrations in, for example, the European Union, flushing out the differences and similarities is crucial [11]. While differences remain, recent research has shown that broader concepts like goals, methods, and structures remain similar across countries within one legal system [12, 13, 14]. These similarities provide a foundational starting point for examining how the concept of Visual Analytics (VA) [15] from computer science can enhance legal research practices. We use the German legal system as an example for our initial analysis, leveraging the expertise of several domain experts available. At the same time, the findings can remain generalizable to at least civil law.
@@
To navigate the landscape of the legal domain, scholars have developed working methods to aid in the interpretation of legal materials, including implicit structures, schemata, and principles [4, p. 450], which are trained and ingrained through years of studies. Visualizations can enhance these methods by improving comprehension and communicating mental models [16, 17]. While the legal domain usually employs visualization for a specific purpose on selected occasions [16, 17], it is an integral part of human-centered and data-driven workflows [15] in computer science. By utilizing computational methods and increasingly available machine-readable legal documents [18], data-driven jurisprudence can efficiently extract knowledge from legal data, complementing traditional doctrinal analysis [19, 20]. Natural Language Processing (NLP) and Artificial Intelligence (AI) have significantly advanced the efficiency and scalability of textual analysis in legal contexts, also driven by Large Language Models (LLMs) [21, 22]. Nevertheless, the transition from linguistic processing to a computational understanding of legal language remains incomplete [23]. Current NLP techniques and AI models frequently lack the nuanced legal comprehension [20] necessary to interpret legal material and infer legal concepts.
@@
At the same time, the digitization of law is fraught with challenges. Substantial portions of legal documents, such as law commentaries, remain restricted by license models and are rarely publicly available. Moreover, some countries lag in legislative efforts to make legal documents available in structured, digital formats [24]. In Germany, for example, less than 1% of associated legal texts are openly available [25]. Consequently, existing open German legal databases (refer to Table 1) offer limited functionality to improve efficiency. Interfaces are primarily keyword-based searches, which often lack context-awareness and fail to account for implicit meanings or relationships [26, 27, 18].
To address these gaps, several commercial solutions have emerged beyond the two primary commercial German databases, beck-online [26] and juris [27], similar to the Anglo-American Westlaw [28] and LexisNexis [29], which primarily feature keyword-search and backlinks. LEX AI offers tools for monitoring regulatory changes and curating personalized feeds with summaries and synopses of regulatory changes supported by AI [30]. LeReTo KG enriches documents with links to related materials [31]. Meanwhile, tools like JURA KI Assistent [32] and Justin Legal [33] assist with tasks like legal drafting and management. The European Union (EU) offers several information retrieval tools for European and national law and jurisdiction, including EUR-Lex [34], N-Lex [35], and the e-Justice Portal [36]. Furthermore, the EU provides LEOS [37], a collaborative online software platform for drafting European legislation with version control. In line with these efforts, the EU advocates for the Interoperable Europe Act [38], which aims to enhance the interoperability of public administrations across its member states by implementing standardized technical solutions for digital services provided to EU citizens. This act also contributes to developing open legal standards, such as the European Legislation Identifier [39].
Despite these advancements, scientific efforts have primarily focused on computational linguistics and text-based analysis [40, 41, 42, 43, 44], leaving a gap in VA tailored to legal workflows. Visualization and VA are critical for supporting knowledge work in fields such as digital humanities and healthcare [45]. Lettieri and Malandrino coined the term Visual Legal Analytics (VLA) and outline challenges in applying VA to legal science, including the large volume of legal texts and the technical unfamiliarity of legal scholars [46]. Additionally, we understand the challenges of VLA as a mismatch between traditional VA models and the workflows of legal scholars.
@@
In legal contexts, traditional VA models are problematic since they often view knowledge as a singular entity resulting from user interactions with visualizations [15, 47]. However, in jurisprudence, it is insufficient to consider legal materials in isolation. Legal reasoning requires the application of domain knowledge to these materials. Hence, VA must distinguish between explicit knowledge, which is found in legal documents and interpretable by machines, and tacit knowledge, which arises from cognitive processing and is accessible only to humans [48, 45].
@@
To address these unique needs of legal scholars, this work identifies previously unexamined opportunities but also challenges in digitalizing jurisprudence, thereby making the following contributions:
@@
  • A case study proposing a VLA design to augment traditional interfaces and to highlight new opportunities in jurisprudence and VA design.
With these contributions, we provide a blueprint for a Visual Legal Analytics design that recognizes tacit domain knowledge as essential for analytical reasoning in jurisprudence and beyond in knowledge work.
@@ 2 Background
Jurisprudence is underpinned by a substantial body of literature outlining methodologies for legal research and interpretation, particularly emphasizing doctrinal analysis [49]. Although jurisprudence traditionally relies on such analysis, legal scholars increasingly recognize the importance of interdisciplinary research for addressing complex legal questions [50]. For example, legal practitioners need to weigh up a conflict of objectives as part of the European General Data Protection Regulation [51] when considering legal aims and technical practicability regarding privacy and trust. Empirical approaches, particularly computational analysis, have become invaluable for supporting evidence-based policy-making and challenging established paradigms in legal research [19]. Siems identifies four distinct types of such interdisciplinary research, differentiating between primarily legal and non-legal questions on one axis and between separated or integrated research methodologies on the other axis [50]. For instance, legal research questions can arise outside the legal domain, such as economics or political sciences, where law is merely one aspect of a broader research question.
2.1 Digital Humanities
Similarly, the Digital Humanities (DH) is an interdisciplinary research field that leverages computational methodologies to transform traditional humanities research paradigms [52, 53]. They facilitate interdisciplinary research by integrating computational methods with qualitative reasoning. While humanities scholars traditionally rely on close reading, i.e., analytical reasoning through textual analysis, which is also common in law, recent developments emphasize the complementary role of distant reading [53]. The distant reading paradigm quantifies and visually abstracts textual data as an integral part of discovery processes, offering new analytical insights [54, 53]. Typical visualization techniques include textual graphs, heat maps, network visualizations, and word clouds [53]. For example, historical textual analyses use visualizations of textual networks or sentiment heatmaps to provide context-sensitive insights into authorship patterns, thematic evolutions, and cultural influences [53]. Oesterling et al. project a high-dimensional feature space of documents employing a metaphor of islands to visualize the relationships of documents in clusters [55]. Jänicke et al. unveil the occurrences of sentences shared between documents of a text corpus using a grid with cells colored like a heat map [56].
To achieve such support, effective visualization design for humanities considers three goals: task specificity, audience requirements, and the underlying source material’s characteristics [52, 57]. In particular, such design communicates its methodological value to the application domain and sets reasonable expectations about its capabilities [52]. Research tasks supported through these visualizations, like tracing thematic developments, interpreting implicit textual meanings, or contextualizing content historically, are similar to interpretive tasks found in jurisprudence, offering a pathway by applying similar techniques for their solution.
2.2 Visual Analytics
Visual Analytics enhances understanding, promotes analytical reasoning, and aids decision-making in large information spaces using automated analysis techniques and interactive visual interfaces [15]. At the beginning of an analysis, an analyst possesses knowledge of the data being analyzed and the problem domain. By interacting with visual interfaces that offer specialized views on the domain-specific data, the analyst guides the exploration to constitute and verify hypotheses [47]. Throughout this process, the analyst acquires further knowledge about the problem domain. However, the process is not systematic and thus requires tracking the provenance of the analysis steps and insights to review the hypotheses [58, 59]. Provenance tracking associates evolving knowledge with each analysis step. In favor of humans’ limited working memory, a mechanism to externalize the evolving knowledge throughout the analysis process is required to establish links with the provenance [58]. Hence, VA is particularly promising for jurisprudence as it integrates human cognitive strengths into computational workflows.
For example, Tian et al. propose LitVis, a VA system for literature exploration and management that supports survey writing [60]. At first, practitioners conduct literature searches to scope their survey. Next, they identify relevant papers and their relationships, leading to topic prioritization and paper classification in the final phase. KAMAS [61] and VIStory [62] adapt similar workflows to interactive storytelling, where researchers analytically reason about trends in scientific communities and the impact of highly-cited publications.
Together, these systems share key abstract tasks of information retrieval, understanding structural relationships, and analytical reasoning with jurisprudence (for details, refer to subsection 3.2). Nevertheless, systematic integration of VA into legal research workflows to support these tasks remains underexplored. This encompasses (1) the dissemination of legal principles and concepts to the general public, (2) the comprehension of legal documents and the formulation of legal arguments, and (3) the assistance of legal professionals in comprehending legislation. Thus, examining how VA can complement existing interdisciplinary methodologies in jurisprudence, particularly through interactive visualization and tacit knowledge, becomes essential.
@@ 2.3 Legal Tech
Legal technology, or legal tech, refers to all technological means that assist users in interacting with the law. This includes the skills and techniques used to operate these means [63]. Importantly, this definition also encompasses information technology, such as visualization and AI. Technological means can be categorized based on their directness and specificity. For instance, while a smartphone can be classified as legal tech, it is not limited to legal purposes and can require considerable effort to be used effectively in that context. In contrast, practical research in legal tech focuses on designing applications specifically for legal needs, which require less input to function effectively. For instance, Landthaler et al. improve semantic search in legal document collections using machine learning (ML) [40]. Similarly, Vianna and Silva De Moura employ topic modeling, an ML technique, to organize Portuguese legal documents [42]. Bachinger et al. abstract specific applications into a workflow that automatically extracts entities from laws using LLMs [44]. To advance these computational methods, Darji et al. present a data set of legal norms annotated by legal experts [64]. Fobbe also provides multiple digitized collections of German legal documents like the German federal law [1]. In the legal community, visual aids such as a tabular representation [16], mind maps, and flowcharts [65, 66, 67], have been explored to communicate legal cases and convey legal concepts. Additionally, other static visualizations enhance the comprehension of law [68, 69, 70, 71, 72]. These include node-link diagrams, glyphs, geographic maps, and hierarchical representations. EUCaseNet and Knowlex complement interactivity that enables information retrieval through visual exploration [73, 74]. In addition, these approaches employ line graphs, heat maps, and tables. La Cava et al. support understanding of structural relationships between legal documents in LawNet-Viz [75]. LegalVis enables analytical reasoning about judicial decisions through AI using bar charts and a text view of legal documents [76]. These visualizations offer various perspectives on legal data. While they aid legal scholars in navigating information spaces, they cannot review the analysis process or facilitate the externalization of knowledge crucial for effective legal reasoning. However, such support is essential to uphold the essence of legal practice. Moreover, merely presenting knowledge or relationships in a different format does not automatically guarantee benefits. The provided capabilities must offer tangible advantages, be accepted, and then be utilized to maximize their transformative potential. To better understand the opportunities that VA can offer for jurisprudence, we must, therefore, discuss legal research practice.

### v3

@@ Abstract
Legal exploration, analysis, and interpretation remain complex and demanding tasks, even for experienced legal scholars, due to the domain-specific language, tacit legal concepts, and intentional ambiguities embedded in legal texts. In related, text-based domains, Visual Analytics (VA) has become an indispensable tool for navigating documents, representing knowledge, and supporting analytical reasoning. However, legal scholarship presents distinct challenges: it requires managing formal legal structure, drawing on tacit domain knowledge, and documenting intricate and accurate reasoning processes – needs that current VA system designs for law fail to address adequately. We identify and describe key challenges and underexplored opportunities in applying VA to law, exploring how these technologies might better serve the legal domain. Interviews with nine legal experts reveal that current legal information retrieval interfaces do not adequately support the navigational complexity of law, often forcing users to rely on internalized legal expertise instead. To address this gap, we identify a three-phase workflow for legal experts, which highlights opportunities for VA to support legal reasoning through knowledge externalization and provenance tracking, leveraging tree-, graph-, and hierarchy-based visualizations. Through this contribution, our work establishes a user-centered VA workflow for the legal domain, recognizing tacit legal knowledge as a critical element of sense-making and insight generation, and situates these contributions within a broader research agenda for VA in law and other text-based disciplines.
Provenance,
@@
[…]
Research Note
{abstractbox}
© The Authors 2025. This is the author’s version of the article that has been published in Springer’s Artificial Intelligence and Law journal. The final version of this article should be cited as: Daniel Fürst, Mennatallah El-Assady, Daniel A. Keim, Maximilian T. Fischer: Challenges and Opportunities for Visual Analytics in Jurisprudence, Artificial Intelligence and Law, Springer. doi: 10.1007/s10506-025-09494-2 (to appear).
@@ 1 Introduction
Legal text data is vast [fobbeCorpusDeutschenBundesrechts2024, fobbeCorpusEntscheidungenBundesverfassungsgerichts2024, fobbeCorpusDrucksachenDeutschen2021] but inherently structured [ruethersRechtstheorieUndJuristische2022]. It combines domain-specific language [ruethersRechtstheorieUndJuristische2022] and precise phraseology with intentional ambiguity at times [ruethersRechtstheorieUndJuristische2022]. This ambiguity allows for statutory interpretation [ruethersRechtstheorieUndJuristische2022], enabling laws to evolve alongside societal values. For instance, in 2018, the German Federal Constitutional Court interpreted Art. 2 Abs. 2 Satz 1 GG to include protection against the dangers of climate change [bundesverfassungsgerichtBundesverfassungsgerichtDecisionsConstitutional2021], where the legal text merely states that “…every person shall have the right to life and physical integrity…” This interpretation underscores that considering legal norms in isolation is insufficient; statutory interpretation also relies on an extensive understanding of fundamental and specific legal principles [ruethersRechtstheorieUndJuristische2022].
@@
These legal principles are formed and conceived within individual jurisdictions and legal systems. Several legal systems co-exist, with common law and civil law being predominant [glennComparativeLegalFamilies2006], while the latter is the most widespread. The Anglo-American common law system relies on judicial precedents and case law, whereas the civil law system, rooted in codes and statutes, lends itself to more structured analysis. These differences are the focus of comparative law studies, reaching back centuries [Schmitthoff.ComparativeLaw.1939, Zimmermann.ComparativeLawHistory.1996]. Seminal work [Fikentscher.MethodenRechtsvergleich.1975, David.MajorLegalSystems.1978] has analyzed comparability of methodologies and approaches, finding that legal practices can differ between jurisdictions and even within legal systems. In particular, with tighter integrations in, for example, the European Union, contrasting is crucial [Schadbach.BenefitsComparativeLaw.1998]. While differences like inductive vs. deductive reasoning, codification vs. precedent referencing, and systematic codes vs. narrative case opinions affect workflows, recent research has shown that broader concepts like goals, methods, and structures remain similar across countries within one legal system [Rossi.Rechtsvergleichung.2017, Reimann.HandbookComparativeLaw.2019, Zweigert.Rechtsvergleichung.2024].
@@
To navigate the landscape of the legal domain, scholars have developed working methods to aid in the interpretation of legal materials, including implicit structures, schemata, and principles [ruethersRechtstheorieUndJuristische2022, p. 450], which are trained and ingrained through years of demanding studies. Visualizations can enhance these methods by improving comprehension and communicating mental models [passalacquaUsingVisualTechniques1997, brunschwigVisualLawVisual2014]. These approaches provide a foundational starting point for examining, with the help of domain experts in law, how Visual Analytics (VA) [keimVisualAnalyticsDefinition2008] can enhance legal research practices. While the legal domain usually employs visualization for a specific purpose on selected occasions [passalacquaUsingVisualTechniques1997, brunschwigVisualLawVisual2014], in Human-Computer-Interaction (HCI) it is an integral part of human-centered and data-driven workflows [keimVisualAnalyticsDefinition2008]. By utilizing computational methods and increasingly available machine-readable legal documents [vogelRichterKeinBot2024], data-driven jurisprudence can efficiently extract knowledge from legal data, complementing traditional doctrinal analysis [lettierinicolaKnowledgeMachineriesIntroducing2019, vaciagogiuseppeOpportunitiesChallengesLegal2019]. Natural Language Processing (NLP) and Artificial Intelligence (AI) have significantly advanced the efficiency and scalability of textual analysis in legal contexts, also driven by Large Language Models (LLMs) [brownLanguageModelsAre2020]. Nevertheless, the transition from linguistic processing to a computational understanding of legal language remains incomplete [baker2018LegalResearch2017]. Current NLP techniques and AI models frequently lack the nuanced legal comprehension [vaciagogiuseppeOpportunitiesChallengesLegal2019] necessary to interpret legal material and infer legal concepts.
@@
At the same time, the digitization of law is fraught with challenges. Substantial portions of legal documents remain restricted by license models and are rarely publicly available. Moreover, some countries lag in legislative efforts to make legal documents available in structured, digital formats [zanderUrteileUnterVerschluss2024]. In Germany, for example, less than 1% of associated legal texts are openly available [ltoTransparenzJustizStagnation]. Consequently, existing open German legal databases (refer to 1) offer limited functionality to improve efficiency. Interfaces are primarily keyword-based searches, which often lack context-awareness and fail to account for implicit meanings or relationships [verlagc.h.beckohgBeckonline, jurisgmbhJuris, vogelRichterKeinBot2024].
@@
Despite these advancements, scientific efforts have primarily focused on computational linguistics and text-based analysis [Landthaler2016ExtendingFT, salaunConditionalAbstractiveSummarization2022, viannaOrganizingPortugueseLegal2022, correiaDynamicTopicModeling2023, bachingerExtractingLegalNorm2024], leaving a gap in VA tailored to legal workflows. Visualization and VA are critical for supporting knowledge work in fields such as digital humanities and healthcare [federicoRoleExplicitKnowledge2017]. We understand the challenges of Visual Legal Analytics (VLA) [lettieriCartographiesLegalWorld2018] as a mismatch between traditional VA models and the workflows of legal scholars. In legal contexts, traditional VA models are problematic since they often view knowledge as a singular entity resulting from user interactions with visualizations [keimVisualAnalyticsDefinition2008, sachaKnowledgeGenerationModel2014]. However, in jurisprudence, it can be insufficient to consider legal materials in isolation, requiring the application of domain knowledge. Hence, VA must distinguish between explicit, machine-interpretable knowledge, found in legal documents, and tacit knowledge, which is domain-specific, personal, and arises only from cognitive processing [wangDefiningApplyingKnowledge2009, federicoRoleExplicitKnowledge2017]. In the legal context, tacit knowledge considers the knowledge about the inherent relationships between the laws and their application, which is often complicated to articulate [ruethersRechtstheorieUndJuristische2022].
@@
To address these unique needs of legal scholars, this work asks the following research question: In what ways can Visual Analytics support legal research by enhancing legal analytical reasoning? It identifies previously unexamined opportunities but also challenges in digitalizing jurisprudence, thereby making the following contributions:
@@
  • A case study proposing a VLA design to augment traditional interfaces, highlighting new opportunities in jurisprudence.
With these contributions, we provide a blueprint for a Visual Analytics design, considering tacit domain knowledge as essential for law applications and beyond.
@@ 2 Background
Jurisprudence is underpinned by methodologies and doctrinal analysis [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024]. Although jurisprudence traditionally relies on such analysis, legal scholars increasingly recognize the importance of interdisciplinary research for addressing complex legal questions [siemsTAXONOMYINTERDISCIPLINARYLEGAL2009]. For example, legal practitioners need to weigh up conflicting objectives as part of the European GDPR [gdpr2016] when considering legal aims and technical practicability regarding privacy and trust.
2.1 Digital-Assisted Law Research
Law research is primarily supported through commercial databases, like beck-online [verlagc.h.beckohgBeckonline] and juris [jurisgmbhJuris] in Germany, similar to the Anglo-American Westlaw [thomsonreutersWestlawClassic] and LexisNexis [lexisnexislegal&professionalLexisNexis]. Those primarily feature keyword-search and backlinks. LEX AI offers tools for monitoring regulatory changes and curating personalized feeds with summaries and synopses of regulatory changes supported by AI [lexaigmbhLEXAI]. [leretokgLeReToLegalResearch] enriches documents with links to related materials [leretokgLeReToLegalResearch]. Meanwhile, tools like JURA KI Assistent [ra-microsoftwareagJURAKIAssistent] and Justin Legal [justinlegaltechgmbhJustinLegalDigitale2023] assist with tasks like legal drafting and management. The European Union (EU) offers several information retrieval tools for European and national law and jurisdiction, including EUR-Lex [publicationsofficeoftheeuropeanunionEURLex], N-Lex [publicationsofficeoftheeuropeanunionNLex], and the e-Justice Portal [directorate-generalforjusticeandconsumersEuropeanEJusticePortal]. Furthermore, the EU provides LEOS [europeancommissionLEOS2015], a collaborative online software platform for drafting European legislation with version control. The EU advocates for the Interoperable Europe Act [theeuropeanparliamentandthecounciloftheeuropeanunionInteroperableEuropeAct2024], which aims to enhance the interoperability of public administrations across its member states by implementing standardized technical solutions provided to EU citizens. This act contributes to developing open legal standards, such as the European Legislation Identifier [publicationsofficeoftheeuropeanunionELI2017].
2.2 Visual Analytics for Law
Close and distant reading [janickeVisualTextAnalysis2017] are well-established concepts in the Digital Humanities (DH). The distant reading paradigm quantifies and visually abstracts textual data as an integral part of discovery processes, offering new analytical insights [morettiGraphsMapsTrees2007, janickeVisualTextAnalysis2017]. At the same time, Visual Analytics enhances understanding, promotes analytical reasoning, and aids decision-making in large information spaces using automated analysis techniques and interactive visual interfaces [keimVisualAnalyticsDefinition2008]. This process is not systematic and thus requires tracking the provenance of the analysis steps and insights to review the hypotheses [shrinivasanSupportingAnalyticalReasoning2008, gotzCharacterizingUsersVisual2008]. In favor of humans’ limited working memory, a mechanism to externalize the evolving knowledge throughout the analysis process is required to establish links with the provenance [shrinivasanSupportingAnalyticalReasoning2008]. Hence, VA is particularly promising for jurisprudence as it integrates human cognitive strengths into computational workflows. For example, [tianLitVisVisualAnalytics2023] propose LitVis for literature exploration and management that supports survey writing [tianLitVisVisualAnalytics2023], conducting search, ranking, prioritization and classification. KAMAS [wagnerKnowledgeassistedVisualMalware2017] and VIStory [zengVIStoryInteractiveStoryboard2021] adapt similar workflows to interactive storytelling, analytically reasoning about trends in scientific communities and the impact of highly-cited publications. Together, these systems share key abstract tasks of information retrieval, understanding structural relationships, and analytical reasoning with jurisprudence (for details, see 3.2). Nevertheless, systematic integration of VA into legal research workflows to support these tasks remains underexplored. This encompasses (1) the dissemination of legal principles to the general public, (2) the comprehension of legal documents and the formulation of arguments, and (3) the assistance of legal professionals in comprehending legislation. Thus, examining how VA can complement existing interdisciplinary methodologies in jurisprudence, particularly through interactive visualization and tacit knowledge, becomes essential.
@@ 2.3 Legal Tech
Legal technology, or legal tech, refers to all technological means that assist users in interacting with the law. This includes the skills and techniques used to operate these means [whalenDefiningLegalTechnology2022]. Importantly, this definition also encompasses information technology, such as visualization and AI. Technological means can be categorized based on their directness and specificity. For instance, while a smartphone can be classified as legal tech, it is not limited to legal purposes and can require considerable effort to be used effectively in that context. In contrast, practical research in legal tech focuses on designing applications specifically for legal needs, which require less input to function effectively. For instance, [Landthaler2016ExtendingFT] improve semantic search in legal document collections using machine learning (ML) [Landthaler2016ExtendingFT]. Similarly, [viannaOrganizingPortugueseLegal2022] employ topic modeling, an ML technique, to organize Portuguese legal documents [viannaOrganizingPortugueseLegal2022]. [bachingerExtractingLegalNorm2024] abstract specific applications into a workflow that automatically extracts entities from laws using LLMs [bachingerExtractingLegalNorm2024]. To advance these computational methods, [darjiDatasetGermanLegal2023] present a dataset of legal norms annotated by legal experts [darjiDatasetGermanLegal2023]. [fobbeCorpusDeutschenBundesrechts2024] also provides multiple digitized collections of German legal documents like the German federal law [fobbeCorpusDeutschenBundesrechts2024]. In the legal community, visual aids such as a tabular representation [passalacquaUsingVisualTechniques1997], mind maps, and flowcharts [khalilQuoVadisVisuelle2014, cantatoreMakingConnectionsIncorporating2016, mclachlanVisualisationLawLegal2021], have been explored to communicate legal cases and convey legal concepts. Additionally, other static visualizations enhance the comprehension of law [merklExplorationLegalText1997, katzMeasuringComplexityLaw2014, gomez2015understanding, ourednikVisualApproachHistory2017, bokwonleeNetworkStructureReveals2018]. These include node-link diagrams, glyphs, geographic maps, and hierarchical representations. EUCaseNet and Knowlex complement interactivity that enables information retrieval through visual exploration [lettieriComputationalApproachExperimental2016, lettieriLegalMacroscopeExperimenting2017]. In addition, these approaches employ line graphs, heat maps, and tables. [lacavaLawNetVizWebbasedSystem2022] support the understanding of structural relationships between legal documents in LawNet-Viz [lacavaLawNetVizWebbasedSystem2022]. LegalVis enables analytical reasoning about judicial decisions through AI using bar charts and a text view of legal documents [resckLegalVisExploringInferring2023]. These visualizations offer various perspectives on legal data. While they aid legal scholars in navigating information spaces, they cannot review the analysis process or facilitate the externalization of knowledge crucial for effective legal reasoning. However, such support is essential to uphold the essence of legal practice. Moreover, merely presenting knowledge or relationships in a different format does not automatically guarantee benefits. The provided capabilities must offer tangible advantages, be accepted, and then be utilized to maximize their transformative potential. To better understand the opportunities that VA can offer for jurisprudence, we must, therefore, discuss legal research practice.